OpenStudy (anonymous):

Simplify: 3i[2i(1+i)-(2-10i)]

OpenStudy (anonymous):

first do the inner bracket 2i(1+i)-(2-10i)=2i+2i^2-2+10i=12i-4 3i(12i-4)=36i^2-12i=-12i-36